Summary Report – November 14th

Crew 215 Sol Summary Report 14-NOV-2019
Sol: 4
Summary Title: Making Progress
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: After ascertaining that Curiosity rover was operational (it was) our approved EVA for this Sol saw us returning to the micrometeorite sample grid in the stream wash at MM5 (518462E 4248949N). There, Guy, Larissa and Steve collected 30 samples before returning to the hab for lunch. 40% of the plots have now been sampled. During this time, gypsum samples collected from the White Moon location (517724E 4254500N) were begun to be cleaned and weighed in preparation for experimentation. This continued through the afternoon where samples from the micrometeorite grid were added to those being processed and packaged. Despite a rehearsal, construction in the hab caused a postponement of the disabled Marstronaut evacuation during the evening engineering EVA and, consequently, a ‘picnic’ dinner was held in the Green Hab.
Look Ahead Plan: Planning an EVA to MM5 (518462E 4248949N) to resume collection of micrometeorite samples and a demonstration of returning a disabled astronaut into the hab during the afternoon engineering EVA that had been postponed from the previous sol. See EVA requests. Continue monitoring of psycho-social stressors.
Anomalies in work: N/A.
Weather: Marginally below freezing overnight, light winds, clouded over during the day.
Crew Physical Status: Healthy
EVA: Micrometeorite sampling EVA. See EVA report
Reports to be file: Sol Summary Report, Operations Report, Science Report, GreenHab Report, EVA Request, EVA Report, Journalist Report, Daily Photos.
Support Requested: N/A

Copyright © The Mars Society. All rights reserved. | Main Site